### v2

@@ 1 Introduction
„Given $\rho>0$ and a nonnegative function $g$, does there exist a point process with density $\rho$ and radial distribution function $g$?“
@@ 2.1 The Kirkwood closure process
Here $j_{\Lambda}^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=j_{\Lambda}^{(n+k)}(x_{1},\dots,x_{n},y_{1},\dots,y_{k})$ for brevity. If the point process $\mathsf{P}$ is stationary then the correlation functions are translationally invariant, and one can write $\rho^{(n)}=\rho^{n}g^{(n)}$ for appropriate functions $g^{(n)}$ depending on $n-1$ variables, where $\rho$ is the so-called intensity or density of the point process. For $n=2$ the function $g=g^{(2)}$ is the so-called radial distribution function.
@@
$1+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(k)}(\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\geq 0$ (9)
@@ 2.2 The Kirkwood-Salsburg operator
$\sup_{n\geq 1}C_{\beta}\,(u)^{n}\|(\boldsymbol{\Pi}\boldsymbol{K}\boldsymbol{\theta})^{(n)}\|_{\infty}$ $\leq C_{\beta}\,(u)^{n}e^{2\beta B}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}|f_{\beta}\,(x_{i_{*}}-y_{j})|C_{\beta}\,(u)^{-n-k+1}\|\boldsymbol{\theta}\|\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
@@ 3 Existence of the Kirkwood closure process
For each $n\in\mathbb{N}$ and $x_{1},\dots,x_{n}\in\Lambda^{n}$ fixed either $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\equiv 0$ for all $z\in B_{z_{0}}$ or $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})/z^{n}$ does not have any zeros in $B_{z_{0}}$.
Proof 3.6.
Let $z>0$ and fix $x_{1},\dots,x_{n}\in\Lambda^{n}$.
  • 1.
    Assume first that $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})=0$. For $n=1$ it follows from $\theta^{(1)}_{\Lambda}(z;x)\geq z/\Xi_{\Lambda}(z)$ that $z=0$ which is a contradiction. Let now $n\geq 2$, since all terms in (24) are nonnegative, this means all terms of the series have to be equal to zero, i.e.
    $\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}=0,\qquad\text{ for all }k\geq 0$
    which in particular implies $H(\{\boldsymbol{{x}}_{n}\})=+\infty$ and thus
    $e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}=0,\qquad\text{ for all }k\geq 0\text{ and all }\boldsymbol{{y}}_{k}\in\Lambda^{k}.$ (29)
    Again using (24) it follows that in this case $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\equiv 0$ for all $z\in B_{z_{0}}$.
  • 2.
    Now assume $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})>0$. Using (24) a straightforward calculation gives
    $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})}{z^{n}}\right)=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{1}{\Xi_{\Lambda}(z)}\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\right)$
    $=\frac{1}{\Xi_{\Lambda}(z)}\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k+1}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k+1}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k+1}-\frac{\Xi_{\Lambda}^{\prime}(z)}{\Xi_{\Lambda}(z)^{2}}\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\int_{\Lambda^{k}}e^{-\beta H(\{\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}\})}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}.$
    It thus follows that
    $z^{n+1}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})}{z^{n}}\right)=\int_{\Lambda}\theta^{(n+1)}_{\Lambda}(z;\boldsymbol{{x}}_{n},x)\mathop{}\!\mathrm{d}x-\theta^{(n)}_{\Lambda}(\boldsymbol{{x}}_{n})\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;x)\mathop{}\!\mathrm{d}x$
    and it can be concluded that
    $\int_{\Lambda}\theta^{(n+1)}_{\Lambda}(z;\boldsymbol{{x}}_{n},x)\mathop{}\!\mathrm{d}x-\theta^{(n)}_{\Lambda}(\boldsymbol{{x}}_{n})\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;x)\mathop{}\!\mathrm{d}x=\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})\,z\,\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\log\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})}{z^{n}}\right).$ (30)
    Since the left-hand side of (30) is analytic as product and sum of analytic functions the same follows for the right-hand side and thus the claim is proved.
Remark 3.7.
Note that $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})=0$ for some $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ also implies $\theta^{(n+k)}_{\Lambda}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=0$ for all $k\geq 1$ and any $\boldsymbol{{y}}_{k}\in\Lambda^{k}$ by (29).
Corollary 3.8.
@@
Proof 3.9.
By Proposition 3.5 the sign of $\theta^{(n)}_{\Lambda}(-z;\boldsymbol{{x}}_{n})$ is determined by the sign of $(-z)^{n}$. This proves the claim.
@@
where $(\theta_{\Lambda}^{(n)}(-z;\cdot))_{n\geq 1}$ is the solution of (20) for $-z$. By Corollary 3.8 $\sigma_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n})\geq 0$ for all $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ and since
@@ 4 The Kirkwood closure process is a Gibbs point process
for any $n\geq 1$ and $F\in L^{1}((\mathbb{R}^{d})^{n})$. From (35) and Corollary 3.8 it also follows that $\iota^{(n)}\geq 0$ for all $n\geq 1$. Furthermore, (35) and Remark 2.3 imply that the solution $\boldsymbol{\theta}$ of (34) also satisfies
@@
=\!= $\int_{\Lambda^{n}}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}F(\boldsymbol{{x}}_{n};\{\boldsymbol{{y}}_{k}\}\cup\{\boldsymbol{{x}}_{n}\})\frac{j_{\Lambda}^{(k+n)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})}j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$
@@
    $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is locally stable, i.e. for any $n\geq 1$ and $x,x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ there holds
    $W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})=\sum_{i=1}^{n}u(x_{i}-x)\geq-2B.$ (49)
@@
From this it also follows that $(-1)^{n}\kappa_{\Lambda}(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ and $(-1)^{n}\kappa(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ for all $\eta\in\Gamma_{0}$, $n\geq 1$, $\boldsymbol{{x}}_{n}\in(\mathbb{R}^{d})^{n}$ and $0<z<z_{0}$.

### v3

@@ 1 Introduction
„Given $\rho>0$ and a nonnegative function $g$, does there exist a point process with density $\rho$ and radial distribution function gg?“
@@ 2.1 The Kirkwood closure process
Here $j_{\Lambda}^{(n+k)}(\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=j_{\Lambda}^{(n+k)}(x_{1},\dots,x_{n},y_{1},\dots,y_{k})$ for brevity. If the point process $\mathsf{P}$ is stationary then the correlation functions are translationally invariant, and one can write $\rho^{(n)}=\rho^{n}g^{(n)}$ for appropriate functions $g^{(n)}$ depending on n1n-1 variables, where $\rho$ is the so-called intensity or density of the point process. For $n=2$ the function $g=g^{(2)}$ is the so-called radial distribution function.
@@
$1+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{\Lambda^{k}}\rho^{(k)}(\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\geq 0.$ (9)
@@ 2.2 The Kirkwood-Salsburg operator
$\sup_{n\geq 1}C_{\beta}\,(u)^{n}\|(\boldsymbol{\Pi}\boldsymbol{K}\boldsymbol{\theta})^{(n)}\|_{\infty}$ $\leq C_{\beta}\,(u)^{n}e^{2\beta B}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}|f_{\beta}\,(x_{i_{*}}-y_{j})|C_{\beta}\,(u)^{-n-k+1}\|\boldsymbol{\theta}\|_{C_{\beta}\,(u)}\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}$
@@
Remark 2.2.
Note that $\theta^{(n)}_{\Lambda}(z;\boldsymbol{{x}}_{n})=0$ for some $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ also implies $\theta^{(n+k)}_{\Lambda}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=0$ for all $k\geq 1$ and any $\boldsymbol{{y}}_{k}\in\Lambda^{k}$ by (24) for $z\geq 0$.
@@ 3 Existence of the Kirkwood closure process
Remark 3.10.
If $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is locally stable, i.e. for any $n\geq 1$ and $x,x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ there holds
$W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})=\sum_{i=1}^{n}u(x_{i}-x)\geq-2B,$ (30)
then Proposition 3.9 follows from the proof of Theorem 3.1 in [6].
The idea of the proof is to approximate the potential $u$ by an appropriate potential $u_{\delta}$ that is locally stable and show that the corresponding solutions of (20) converge for $\delta\to 0$.
Proof 3.11.
Let $z\in(0,z_{0})$ be given choose $\delta>0$ such that $z\in(0,z_{\delta})$ where
$z_{\delta}:=\left(e^{2\beta B+1}\exp\left(\delta/C_{\beta}\,(u)\right)C_{\beta}\,(u))\right)^{-1}$
and define
$u_{\delta}:=u+\infty\mathds{1}_{|x|<r_{0}}.$ (31)
Here $r_{0}=r_{0}(\delta)>0$ is chosen such that
$\int_{\mathbb{R}^{d}}\left|f_{\beta}\,^{\delta}(x)\right|\mathop{}\!\mathrm{d}x\leq C_{\beta}\,(u)+\delta$
where $f_{\beta}\,^{\delta}:=e^{-\beta u_{\delta}(\cdot)}-1$. In particular, since $u_{\delta}\geq u$ one can use the same stability constant for $u_{\delta}$ as for $u$. Furthermore, if (16) holds for $u$ it also holds for $u_{\delta}$ and thus the definition of $\mathbf{\Pi}$ does not need to be changed. One can now define $\boldsymbol{K}_{\delta}$ as in (18) and (19) with $u_{\delta}$ in place of $u$ and gets that the corresponding version of (20) has a unique solution $\boldsymbol{\theta}_{\Lambda,\delta}$ for $|z|<z_{\delta}$ since $\|\boldsymbol{\Pi}\boldsymbol{K}_{\delta}\|_{E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}}\leq e^{2\beta B+1}\exp\left(\delta/C_{\beta}\,(u)\right)C_{\beta}\,(u)$. From (22) it follows that
$\|\boldsymbol{\theta}_{\Lambda,\delta}\|_{C_{\beta}\,(u)}\leq|z|C_{\beta}\,(u)\sum_{k=0}^{\infty}|z|^{k}\|\boldsymbol{\chi}_{\Lambda}\boldsymbol{\Pi}\boldsymbol{K}\|_{E_{C_{\beta}\,(u)}\to E_{C_{\beta}\,(u)}}^{k}.=\frac{zC_{\beta}\,(u)}{1-|z|e^{2\beta B+1}\exp\left(\delta/C_{\beta}\,(u)\right)C_{\beta}\,(u)}$ (32)
and it follows that the sequence $(\boldsymbol{\theta}_{\Lambda,\delta})_{\delta>0}$ has a subsequence for $\delta\to 0$ such that for every $n\geq 0$ and $F_{n+1}\in L^{1}((\mathbb{R}^{d})^{n+1})$ there holds
$\lim_{\delta\to 0}\int_{(\mathbb{R}^{d})^{n+1}}F_{n+1}(x,\boldsymbol{{x}}_{n})\theta_{\Lambda,\delta}^{(n+1)}(x,\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}(x,\boldsymbol{{x}}_{n})=\int_{(\mathbb{R}^{d})^{n}}F_{n+1}(x,\boldsymbol{{x}}_{n})\theta_{\Lambda,*}^{(n+1)}(x,\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}(x,\boldsymbol{{x}}_{n})$ (33)
for some $\boldsymbol{\theta}_{\Lambda,*}=(\theta^{(n)}_{\Lambda,*})_{n\geq 1}$. Since $\boldsymbol{\theta}_{\Lambda,\delta}$ satisfies the Kirkwood-Salsburg equations for the modified potential $u_{\delta}$ one can conclude that
$\int_{(\mathbb{R}^{d})^{n+1}}F_{n+1}(x,\boldsymbol{{x}}_{n})\theta_{\Lambda,\delta}^{(n+1)}(x,\boldsymbol{{x}}_{n})\mathop{}\!\mathrm{d}(x,\boldsymbol{{x}}_{n})$
$=\int_{(\mathbb{R}^{d})^{n}}F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W_{\delta}(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\left(\theta_{\Lambda,\delta}^{(n)}(\boldsymbol{{x}}_{n})+\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\int_{(\mathbb{R}^{d})^{k}}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})\theta_{\Lambda,\delta}^{(n+k)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\right).$
again with the convention $\theta_{\Lambda,\delta}^{(0)}:=1$. Define
$\widetilde{F}^{\delta}_{n+1+k}(x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k}):=F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W_{\delta}(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})$ (34)
then there holds
$\lim_{\delta\to 0}\widetilde{F}^{\delta}_{n+1+k}(x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})$
almost everywhere and furthermore by (16) there holds
$\left|F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W_{\delta}(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})\right|\leq\left|F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{2\beta B}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})\right|$ (35)
using dominated convergence it can thus be concluded that
$\lim_{\delta\to 0}\int_{(\mathbb{R}^{d})^{n+1+k}}\left|\widetilde{F}^{\delta}_{n+1+k}(x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})-F_{n+1}(x,\boldsymbol{{x}}_{n})ze^{-\beta W(\{x\}\mid\{\boldsymbol{{x}}_{n}\})}\prod_{j=1}^{k}f_{\beta}\,^{\delta}(x-y_{j})\right|\mathop{}\!\mathrm{d}(x,\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})=0.$ (36)
The $L^{1}$ convergence of the $\widetilde{F}^{\delta}_{n+1+k}$ together with (33) then shows that the limit $\boldsymbol{\theta}_{\Lambda,*}$ satisfies the Kirkwood-Salsburg equations for the original potential $u$ and satisfies (29) since every $\boldsymbol{\theta}_{\Lambda,\delta}$ satisfies (29) by Remark 3.10. Since this solution is unique the Proposition is proved.
@@
where $(\theta_{\Lambda}^{(n)}(-z;\cdot))_{n\geq 1}$ is the solution of (20) for z-z. By Proposition 3.9 $\sigma_{\Lambda}^{(n)}(\boldsymbol{{x}}_{n})\geq 0$ for all $\boldsymbol{{x}}_{n}\in\Lambda^{n}$ and since
@@ 4 The Kirkwood closure process is a Gibbs point process
for any $n\geq 1$ and $F\in L^{1}((\mathbb{R}^{d})^{n})$. From (40) and Corollary LABEL:cor:pos it also follows that $\iota^{(n)}\geq 0$ for all $n\geq 1$. Furthermore, (40) and Remark 2.4 imply that the solution $\boldsymbol{\theta}$ of (39) also satisfies
@@
$=\!$ $\int_{\Lambda^{n}}\sum_{k=0}^{\infty}\frac{1}{k!}\int_{\Lambda^{k}}F(\boldsymbol{{x}}_{n};\{\boldsymbol{{y}}_{k}\}\cup\{\boldsymbol{{x}}_{n}\})\frac{j_{\Lambda}^{(k+n)}(z;\boldsymbol{{x}}_{n},\boldsymbol{{y}}_{k})}{j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})}j_{\Lambda}^{(k)}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}\boldsymbol{{y}}_{k}\mathop{}\!\mathrm{d}\boldsymbol{{x}}_{n}.$
@@
    $u\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is locally stable.
@@
Lastly, one can conclude that $(-1)^{n}\kappa_{\Lambda}(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ and $(-1)^{n}\kappa(-z;\boldsymbol{{x}}_{n};\eta)\geq 0$ for all $\eta\in\Gamma_{0}$, $n\geq 1$, $\boldsymbol{{x}}_{n}\in(\mathbb{R}^{d})^{n}$ and $0<z<z_{0}$.
Proof 4.21.
Let $z\in(0,z_{0})$ and fix $x_{1},\dots,x_{n}\in\Lambda^{n}$.
  • 1.
    Assume first that $\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})=0$. For $k=1$ it follows from $\theta^{(1)}_{\Lambda}(z;x)\geq z/\Xi_{\Lambda}(z)$ that $z=0$ which is a contradiction. Let now $k\geq 2$, since all terms in (24) are nonnegative, this means all terms of the series have to be equal to zero, i.e.
    $\int_{\Lambda^{l}}e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l}=0,\qquad\text{ for all }l\geq 0$
    which in particular implies $H(\{\boldsymbol{{y}}_{l}\})=+\infty$ and thus
    $e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}=0,\qquad\text{ for all }l\geq 0\text{ and all }{\boldsymbol{{w}}}_{l}\in\Lambda^{l}.$ (64)
    Again using (24) it follows that in this case $\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})\equiv 0$ for all $z\in B_{z_{0}}$.
  • 2.
    Now assume $\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})>0$. Using (24) a straightforward calculation gives
    $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(n)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}{z^{k}}\right)=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{1}{\Xi_{\Lambda}(z)}\sum_{l=0}^{\infty}\frac{z^{l}}{l!}\int_{\Lambda^{l}}e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l}\right)$
    $=\frac{1}{\Xi_{\Lambda}(z)}\sum_{l=0}^{\infty}\frac{z^{l}}{l!}\int_{\Lambda^{l+1}}e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l+1}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l+1}-\frac{\Xi_{\Lambda}^{\prime}(z)}{\Xi_{\Lambda}(z)^{2}}\sum_{l=0}^{\infty}\frac{z^{l}}{l!}\int_{\Lambda^{l}}e^{-\beta H(\{\boldsymbol{{y}}_{k},{\boldsymbol{{w}}}_{l}\})}\mathop{}\!\mathrm{d}\boldsymbol{{w}}_{l}.$
    It thus follows that
    $z^{k+1}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\left(\frac{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}{z^{k}}\right)=\int_{\Lambda}\theta^{(k+1)}_{\Lambda}(z;\boldsymbol{{y}}_{k})\mathop{}\!\mathrm{d}y-\theta^{(k)}_{\Lambda}(\boldsymbol{{y}}_{k})\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;y)\mathop{}\!\mathrm{d}y$
    and it can be concluded that
    $\frac{1}{\theta^{(y)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}\int_{\Lambda}\theta^{(k+1)}_{\Lambda}(z;\boldsymbol{{y}}_{k},y)\mathop{}\!\mathrm{d}y-\int_{\Lambda}\theta^{(1)}_{\Lambda}(z;y)\mathop{}\!\mathrm{d}y=\,z\,\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}z}\log\left(\frac{\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})}{z^{k}}\right).$ (65)
    Since the left-hand side of (65) is analytic as a solution of (57) the same follows for the right-hand side and thus the the sign of $\theta^{(k)}_{\Lambda}(z;\boldsymbol{{y}}_{k})\neq 0$ is determined by the sign of $z$.